OpenStudy (ggdvaaa):

Pleasseee helppp ) Using the GCF you found in Part B, rewrite 72 + 81 as two factors. One factor is the GCF and the other is the sum of two numbers that do not have a common factor.

OpenStudy (ggdvaaa):

The GCF of 72 and 81 is 9

Directrix (directrix):

72 + 81 = 9* ( 8 + ? )

Directrix (directrix):

Distributive Property

OpenStudy (ggdvaaa):

So 9 times 8?

Directrix (directrix):

@ggdvaaa 9 times what is 81

Directrix (directrix):

Try again.

OpenStudy (ggdvaaa):

9 times 9

Directrix (directrix):

72 + 81 = 9* ( 8 + 9 ) = 9*17

Directrix (directrix):

This is two factors: 9* ( 8 + 9 ) But, I think the question is asking that the two numbers be added to give: 9*17 as the two factors.

Directrix (directrix):

@ggdvaaa

OpenStudy (ggdvaaa):

was that the answer @Directrix

OpenStudy (ggdvaaa):

?

Directrix (directrix):

>> 9*17 is what we got.

OpenStudy (ggdvaaa):

I put like that? 9*17 exactly?

OpenStudy (ggdvaaa):

and thank youu

OpenStudy (ggdvaaa):

@Directrix

Directrix (directrix):

I can't you your answer format box on the homework computer, so I don't know how to tell you to enter it. 9 times 17 What do you see on the answer form for the computer?